I dont understand how to plug it into the formula, please help.
A basketball was thrown from one end of a court to the other. It had an initial upward velocity of 48 feet per second. How long did it take for the ball to hit the ground? Use the formula h = vt – 16t2.

Answer by edjones(8007)

v=velocity=48, h=height=0, t=time
h = vt – 16t^2
48t-16t^2=0
16t(3-t)=0
-t=-3
t=3 sec
.
Ed